OpenStudy (anonymous):

5. What is the equation of the line that contains the points with (x, y) coordinates (–3, 7) and (5, –1)? A. y = 3x – 2 B. y = x + 10 C. y = –x + 4 D. y=−1/3x+8 E. y=−3/2x+114

OpenStudy (anonymous):

and how did you get the answer

OpenStudy (anonymous):

C. y=-x+4 leave the choices find equation of line by two point form and u will see it is y=-x+4

OpenStudy (aroub):

first u have to find the slope (do u know how to find the slope??) -1-7/5+3 -8/8 -1 second u choose any of the points u have (5,-1) ,(x,y) then u substitute theses points in the form y=mx+b ( the 5 by x and the y -1 ) to find the y-intercept which is b -1=-1(5)+b -1=-5+b -1+5=b 4=b so we got the y-intercept and we put it again in the form y=mx+b y=-x+4

OpenStudy (aroub):

did u understand??

OpenStudy (anonymous):

thanks a lot :)

OpenStudy (aroub):

well aaannnyyttiimmeee!!!!! :D :D